OpenStudy (anonymous):

given the values of x and y, find an equation of variation in which y varies directly as x. y=400, when x=110

OpenStudy (anonymous):

y = k x where k is constant 400 = k 110 k = 40/11 so, y = 40/11 x 11y=40x

OpenStudy (anonymous):

i have x=3.64y y=3.64x y=0.28x x=290y

OpenStudy (across):

\[\frac{40}{11}\approx3.64\]

OpenStudy (anonymous):

across...is right

OpenStudy (anonymous):

so y=3.64x?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

hey across

OpenStudy (anonymous):

on the slope question tht answer seems too easy?